# Unified Parameter-Efficient Unlearning for LLMs

### v1

@@ E.1 Errors analysis for Taylor expansion approximation
Since $\hat{\Theta}$ is an optimal solution to the loss function defined in Equation 3 in the submitted manuscript, we have $\nabla_{\Theta}R(\mathcal{Z};\hat{\Theta})=0$. Therefore,
@@
Since $\hat{\Theta}$ is an optimal solution to the loss function defined in Equation 3 in the submitted manuscript, $H_{\hat{\Theta}}$ is positive semidefinite. Therefore, the assumption that $H_{\hat{\Theta}}$ is invertiable implies that $H_{\hat{\Theta}}$ is positive definte. Therefore, we know that
@@
E.2 Errors analysis for our proposed algorithm
For our proposed Algorithm, the estimation errors analysis is as follows. For a given (approximate) solution $\widetilde{\Delta}$ to the Equation 12 in our manuscript, the error is defined as
@@ E.2 Errors analysis for our proposed algorithm
where the function $F(\cdot)$ is defined in Equation 14 in the submitted manuscript. Therefore, the theoretical analysis of $err(\widetilde{\Delta})$ is equivalent to the error analysis of $\|\nabla F(\Delta_{t})\|$ for the sequence $\{\Delta_{t}\}_{t\geq 1}$ generated by the optimization algorithm for solving the problem Equation 9, Equation 10, and Equation 11 in the submitted manuscript.
@@
Appendix F Discussion about the efficiency of LLMEraser
@@
Specifically, LLMEraser formulates the parameter updates as an inverse Hessian-vector product (Equation 9, Equation 10, and Equation 11 in our manuscript). Importantly, although the inverse Hessian appears in the formulation, it does not require explicit computation or inversion of the Hessian matrix. Directly calculating the inverse Hessian-vector product has a time complexity of $O(p^{3})$ and a space complexity of $O(p^{2})$, as the Hessian matrix needs to be stored—making it highly memory-intensive.
Our method transforms the computation of the inverse Hessian-vector product into the problem of solving for the Hessian-vector product, enabling efficient resolution through mini-batch algorithms. The Hessian-vector product, if computed directly via the full Hessian matrix multiplication, would have a time and space complexity of $O(p^{2})$. However, using HVP (Hessian-free methods), we avoid the explicit computation and storage of the Hessian matrix, reducing both time and space complexity to $O(p)$ (Pearlmutter, 1994). By further leveraging mini-batch optimization for Equation 12 in the manuscript, LLMEraser achieves a space complexity of $O(p)$, ensuring its scalability.
@@
Table 9: Experimental results on the QM task for different LoRA ranks (8, 16, 32), using LLaRA as the LLM4Rec model on the LastFM dataset, where 10% of users have items replaced with noisy interactions. ”Corrupted” refers to the model trained with the noisy data.
@@
The GPU utilization of SISA is identical to that of Retrain because SISA (Kwak et al., 2017) effectively requires retraining all parameters (We report the memory usage required to train a single shard). Similarly, fine-tuning-based methods such as gradient descent also necessitates updating all parameters. The backbone of the LLM we used is LLaMA2-7B (Touvron et al., 2023b). Our method is highly memory-efficient and can be executed on a single A40 GPU.
@@
The runtime results for LoRA with ranks 8, 16, and 32 on the LastFM dataset are shown in Table 10. The evaluation is measured in seconds (s).
@@
Appendix G Related Work
G.1 Large Language Models
@@
LLMs have found extensive applications in various downstream tasks, demonstrating their versatility across domains such as natural language processing, information retrieval, and knowledge graph augmentation (Hu et al., 2024a; Zhang et al., 2024a; Xu et al., 2024b). For instance, LLMs are employed to enhance the accuracy of query-based systems by leveraging their ability to understand and generate contextually relevant responses, improving user experience in search applications (Liu et al., 2024b; Shang & Huang, 2024). Additionally, they are utilized in graph analytics, enabling complex reasoning tasks and facilitating the extraction of insights from structured data (Chen et al., 2023; Xu et al., 2024a). The adaptability of LLMs through prompt engineering further supports their deployment in specific use cases, allowing for tailored outputs that meet diverse requirements (Arawjo et al., 2024; Cain, 2024).
G.2 Large Language Models Unlearning

### v2

@@ D.1 Errors Analysis for Taylor Expansion Approximation
Since $\hat{\Theta}$ is an optimal solution to the loss function defined in Equation 3, we have $\nabla_{\Theta}R(\mathcal{Z};\hat{\Theta})=0$. Therefore,
@@
Since $\hat{\Theta}$ is an optimal solution to the loss function defined in Equation 3, $H_{\hat{\Theta}}$ is positive semidefinite. Therefore, the assumption that $H_{\hat{\Theta}}$ is invertiable implies that $H_{\hat{\Theta}}$ is positive definte. Therefore, we know that
@@
D.2 Errors analysis for our proposed algorithm
For our proposed Algorithm, the estimation errors analysis is as follows. For a given (approximate) solution $\widetilde{\Delta}$ to the Equation 12, the error is defined as
@@ D.2 Errors analysis for our proposed algorithm
where the function $F(\cdot)$ is defined in Equation 14. Therefore, the theoretical analysis of $err(\widetilde{\Delta})$ is equivalent to the error analysis of $\|\nabla F(\Delta_{t})\|$ for the sequence $\{\Delta_{t}\}_{t\geq 1}$ generated by the optimization algorithm for solving the problem Equation 9, Equation 10, and Equation 11.
@@
Appendix E Discussion About the Efficiency of LLMEraser
@@
Specifically, LLMEraser formulates the parameter updates as an inverse Hessian-vector-product (Equation 9, Equation 10, and Equation 11). Importantly, although the inverse Hessian appears in the formulation, it does not require explicit computation or inversion of the Hessian matrix. Directly calculating the inverse Hessian-vector-product has a time complexity of $O(p^{3})$ and a space complexity of $O(p^{2})$, as the Hessian matrix needs to be stored—making it highly memory-intensive.
Our method transforms the computation of the inverse Hessian-vector-product into the problem of solving for the Hessian-vector-product, enabling efficient resolution through mini-batch algorithms. The Hessian-vector-product, if computed directly via the full Hessian matrix multiplication, would have a time and space complexity of $O(p^{2})$. However, using HVP (Hessian-free methods), we avoid the explicit computation and storage of the Hessian matrix, reducing both time and space complexity to $O(p)$ (Pearlmutter, 1994). By further leveraging mini-batch optimization for Equation 12, LLMEraser achieves a space complexity of $O(p)$, ensuring its scalability.
@@
Table 8: Experimental results on the QM task for different LoRA ranks (8, 16, 32), using LLaRA as the LLM4Rec model on the LastFM dataset, where 10% of users have items replaced with noisy interactions. “Corrupted” refers to the model trained with the noisy data.
@@
The GPU utilization of SISA is identical to that of Retrain because SISA (Kwak et al., 2017) effectively requires retraining all parameters (We report the memory usage required to train a single shard). Similarly, fine-tuning-based methods such as gradient descent also necessitates updating all parameters. The backbone of the LLM we used is LLaMA2-7B (Touvron et al., 2023b).
@@
The runtime results for LoRA with ranks 8, 16, and 32 on the LastFM dataset are shown in Table 9. The evaluation is measured in seconds.
@@
Appendix F Related Work
F.1 Large Language Models
@@
LLMs have found extensive applications in various downstream tasks (Fang et al., 2024b; Hu et al., 2024a; Wu et al., 2024b), demonstrating their versatility across domains such as natural language processing, information retrieval, and knowledge graph augmentation (Zhang et al., 2024a; Xu et al., 2024b; Fang et al., 2024a; Sheng et al., 2024). For instance, LLMs are employed to enhance the accuracy of query-based systems by leveraging their ability to understand and generate contextually relevant responses, improving user experience in search applications (Liu et al., 2024b; Shang & Huang, 2024). Additionally, they are utilized in graph analytics, enabling complex reasoning tasks and facilitating the extraction of insights from structured data (Chen et al., 2023; Xu et al., 2024a). The adaptability of LLMs through prompt engineering further supports their deployment in specific use cases, allowing for tailored outputs that meet diverse requirements (Arawjo et al., 2024; Cain, 2024).
In a similar vein, LLMs are increasingly being integrated into recommendation systems, building on their capabilities in natural language processing and understanding user preferences. Traditional recommendation systems often rely on collaborative filtering (Misztal-Radecka & Indurkhya, 2020; Wu et al., 2024a), content-based approaches (Pazzani & Billsus, 2007; Wu et al., 2022), or hybrid models (Burke, 2002). Recent advances, including Reinforced Prompt Personalization (Mao et al., 2024; Xin et al., 2022), and the incorporation of LLMs into recommendation systems via tool learning (Zhao et al., 2024; Dehbozorgi et al., 2024) or fine-tuning with recommendation-specific data (Kong et al., 2024; Chen et al., 2024), have significantly improved personalization. These methods enable LLMs to better capture user preferences and context (Lyu et al., 2024; Hu et al., 2024b), ultimately enhancing the accuracy and relevance of recommendations.
F.2 Large Language Models Unlearning
@@
Appendix G More examples of various unlearning tasks
Figure 5: Response Correction Case Study.